# SlideChain: Semantic Provenance for Lecture Understanding via Blockchain Registration

### v1

@@ 1 Introduction
Multimodal vision–language models (VLMs)have rapidly advanced in their ability and are increasingly being used to interpret complex educational materials, including scientific figures, mathematical formulas, imaging geometries, and lecture slides. Recent surveys and benchmarks highlight progress in multimodal pretraining and downstream task performance [undef], while large instruction-tuned VLMs such as LLaVA, Qwen-VL, and InternVL demonstrate impressive promises in captioning, question answering, knowledge extraction, and diagram interpretation. These capabilities position VLMs as potential tools for augmenting STEM instruction, enhancing lecture understanding, and enabling intelligent tutoring. Yet despite their rapid advancement, the semantic reliability of modern VLMs remains poorly understood, particularly in a field-specific fashion, such as for biomedical imaging education.
@@
  • Multimodel semantic extraction pipeline. We develop a unified extraction framework that produces structured concepts, relational triples, and metadata for each slide in the SlideChain Slides Dataset using four state-of-the-art VLMs.
@@ 2.1 Multimodal Slide Understanding and VLM Limitations
Recent progress in large vision–language models (VLMs) has dramatically expanded the ability of multimodal systems to interpret complex visual and textual inputs, including diagrams, figures, and structured document layouts. Models such as LLaVA [undeff], Qwen-VL [undefg], and InternVL3 [undefh] demonstrate strong capabilities in image-grounded reasoning, caption generation, and instruction-following tasks across diverse real-world settings. Despite these advances, their behavior on structured educational slides—where semantics combine visual layout, mathematical notation, and dense conceptual links—remains insufficiently studied.
Slide understanding has emerged as a distinct research area within multimodal AI. Prior works introduced datasets focusing on text detection, segmentation, or video–slide alignment, including SlideVQA [undefi], educational multimodality datasets such as LPM and M3AV, and lecture-driven multimodal corpora such as the ICCV Lecture Presentations dataset [undefj]. More recently, Zhang et al. proposed the LecSlides-370K dataset [undefk], a large-scale benchmark emphasizing global slide comprehension and question answering across diverse academic fields. These efforts highlight the growing importance of robust slide understanding but also reveal the challenges inherent to multimodal educational materials.
@@ 2.3 Blockchain for Scientific Integrity
In scientific computing, SciChain [undefv] demonstrated the feasibility of deploying blockchain-based provenance in high-performance computing (HPC) environments. The system introduced new consensus protocols tailored to resource-constrained scientific workflows and showed that blockchain can serve as a practical substrate for reproducible computational experiments. Complementing these system-level contributions, Akbarfam and Maleki’s SoK survey [undefw] provides a comprehensive taxonomy of blockchain-for-provenance approaches, highlighting key design choices, security properties, and application domains.
@@ 3.1 Pipeline Summary
The SlideChain pipeline begins with the ingestion of high-resolution slide images and accompanying transcript snippets from the SlideChain Slides Dataset, a curated collection of 1,117 educational slides from a medical imaging course. Each slide is processed independently through four state-of-the-art vision–language models (VLMs), each configured for multimodal concept and relational triple extraction. The resulting outputs differ widely in structure and completeness, necessitating a normalization stage.
@@ 4 Dataset and Models
This work uses the SlideChain Slides Dataset as the semantic substrate for evaluating multimodel extraction, cross-model variability, and blockchain-backed provenance. The dataset consists of educational slides from a full-semester medical imaging course, and four state-of-the-art vision–language models (VLMs) are applied to every slide to generate complementary concept and relation extractions that form the basis of our semantic and provenance analyses.
@@ 5 Semantic Analysis Methods
5.1 Concept and Triple Extraction
@@ 8 Results
This section presents the empirical evaluation of SlideChain across three complementary dimensions. First, we analyze the semantic behavior and variability of four vision–language models (VLMs) on 1,117 educational slides, including disagreement, cross-model similarity, lecture-level patterns, and stability characteristics. Second, we perform a lightweight baseline comparison between single-model and multi-model semantic provenance to quantify how much semantic content is missed when relying on a single VLM. Third, we evaluate the performance, scalability, and reliability of the blockchain-backed provenance layer, including gas consumption, throughput, cost projections, integrity guarantees, tamper detection, and reproducibility across independent extraction runs. All results are reported using the full SlideChain Slides Dataset.
@@ 8.3.3 Cost Analysis
On an Ethereum-like L1 environment, per-slide cost ranges from \$0.69 to \$1.23, totaling approximately \$780 for the 1,117-slide dataset. While L1 deployment remains relatively expensive, Layer-2 (L2) ecosystems such as Optimism- or Polygon-like rollups reduce these costs by one to two orders of magnitude, making continuous provenance registration feasible for real-world educational pipelines.
@@ 9 Discussion
This work presents SlideChain, a provenance infrastructure that couples multimodal semantic extraction with blockchain-backed verification to support auditability and reproducibility in educational AI systems. Rather than proposing improvements to semantic correctness itself, SlideChain focuses on making the outputs of semantic extraction verifiable, traceable, and comparable over time. Across semantic analysis, blockchain evaluation, and reproducibility experiments, our results highlight both the opportunities and the limitations of current vision–language models (VLMs) when applied to scientific and engineering lecture material, and motivate provenance as a necessary system-level capability.
@@ 10 Conclusion
This paper introduced SlideChain, a blockchain-backed provenance framework designed to ensure the integrity, auditability, and long-term reproducibility of multimodal semantic extraction pipelines for educational content. Using the SlideChain Slides Dataset of 1,117 medical imaging lecture slides and four state-of-the-art vision–language models (VLMs), we presented, to the best of our knowledge, the first empirical study of semantic variability in real-world STEM lecture material. Our results reveal substantial cross-model divergence in both concept extraction and relational grounding, underscoring that modern VLMs—despite their impressive generative capabilities—do not yet provide stable or authoritative interpretations of scientific slides. SlideChain addresses this challenge by anchoring cryptographic commitments of slide-level semantic records to an Ethereum-compatible blockchain. While simple hashing or local logging can detect accidental file changes, they cannot provide immutable timestamps, third-party verifiability, or protection against silent overwriting. In contrast, blockchain-backed commitments establish an append-only, tamper-evident ledger that decouples trust from any single system owner and enables independent auditing of semantic interpretations over time. By storing only compact Keccak-256 hashes on-chain and retaining full semantic records off-chain, SlideChain achieves this integrity guarantee without sacrificing flexibility or scalability. Our evaluations demonstrate that the system exhibits predictable, constant gas usage, near-linear scalability, and stable throughput of approximately one slide per second under batch registration. Tamper-detection experiments achieved 100% accuracy, and dual-run reproducibility tests produced perfect Jaccard similarity across all 4,468 model–slide combinations, confirming deterministic behavior under controlled execution settings. Taken together, these results establish SlideChain as a practical and extensible provenance infrastructure for multimodal educational AI. Rather than attempting to resolve semantic correctness, SlideChain provides a verifiable semantic baseline against which model behavior, drift, and disagreement can be transparently audited. As AI-generated interpretations become increasingly embedded in STEM pedagogy and digital learning ecosystems, such provenance guarantees will be essential for maintaining trust, accountability, and long-term reproducibility.

### v2

@@ 1 Introduction
Multimodal VLMs have rapidly advanced in their ability and are increasingly being used to interpret complex educational materials, including scientific figures, mathematical formulas, imaging geometries, and lecture slides. Recent surveys and benchmarks highlight progress in multimodal pretraining and downstream task performance [undef], while large instruction-tuned VLMs such as LLaVA, Qwen-VL, and InternVL demonstrate impressive promises in captioning, question answering, knowledge extraction, and diagram interpretation. These capabilities position VLMs as potential tools for augmenting STEM instruction, enhancing lecture understanding, and enabling intelligent tutoring. Yet despite their rapid advancement, the semantic reliability of modern VLMs remains poorly understood, particularly in a field-specific fashion, such as for biomedical imaging education.
@@
  • Multimodal semantic extraction pipeline. We develop a unified extraction framework that produces structured concepts, relational triples, and metadata for each slide in the SlideChain Slides Dataset using four state-of-the-art VLMs.
@@ 2.1 Multimodal Slide Understanding and VLM Limitations
Recent progress in large vision–language models has dramatically expanded the ability of multimodal systems to interpret complex visual and textual inputs, including diagrams, figures, and structured document layouts. Models such as LLaVA [undeff], Qwen-VL [undefg], and InternVL3 [undefh] demonstrate strong capabilities in image-grounded reasoning, caption generation, and instruction-following tasks across diverse real-world settings. Despite these advances, their behavior on structured educational slides—where semantics combine visual layout, mathematical notation, and dense conceptual links—remains insufficiently studied.
Slide understanding has emerged as a distinct research area within multimodal AI. Prior works introduced datasets focusing on text detection, segmentation, or video–slide alignment, including SlideVQA [undefi], educational multimodality datasets such as Lecture Presentation Modality and Multimodal Multiview Audio-Visual, and lecture-driven multimodal corpora such as the ICCV Lecture Presentations dataset [undefj]. More recently, Zhang et al. proposed the LecSlides-370K dataset [undefk], a large-scale benchmark emphasizing global slide comprehension and question answering across diverse academic fields. These efforts highlight the growing importance of robust slide understanding but also reveal the challenges inherent to multimodal educational materials.
@@ 2.3 Blockchain for Scientific Integrity
In scientific computing, SciChain [undefv] demonstrated the feasibility of deploying blockchain-based provenance in high-performance computing (HPC) environments. The system introduced new consensus protocols tailored to resource-constrained scientific workflows and showed that blockchain can serve as a practical substrate for reproducible computational experiments. Complementing these system-level contributions, Akbarfam and Maleki’s Systematization of Knowledge survey [undefw] provides a comprehensive taxonomy of blockchain-for-provenance approaches, highlighting key design choices, security properties, and application domains.
@@ 3.1 Pipeline Summary
The SlideChain pipeline begins with the ingestion of high-resolution slide images and accompanying transcript snippets from the SlideChain Slides Dataset, a curated collection of 1,117 educational slides from a medical imaging course. Each slide is processed independently through four state-of-the-art VLMs, each configured for multimodal concept and relational triple extraction. The resulting outputs differ widely in structure and completeness, necessitating a normalization stage.
@@ 4 Dataset and Models
This work uses the SlideChain Slides Dataset as the semantic substrate for evaluating multimodel extraction, cross-model variability, and blockchain-backed provenance. The dataset consists of educational slides from a full-semester medical imaging course, and four state-of-the-art VLMs are applied to every slide to generate complementary concept and relation extractions that form the basis of our semantic and provenance analyses.
@@ 5 Semantic Analysis Methods
5.1 Normalization of Concepts and Triples
@@ 8 Results
This section presents the empirical evaluation of SlideChain across three complementary dimensions. First, we analyze the semantic behavior and variability of four VLMs on 1,117 educational slides, including disagreement, cross-model similarity, lecture-level patterns, and stability characteristics. Second, we perform a lightweight baseline comparison between single-model and multi-model semantic provenance to quantify how much semantic content is missed when relying on a single VLM. Third, we evaluate the performance, scalability, and reliability of the blockchain-backed provenance layer, including gas consumption, throughput, cost projections, integrity guarantees, tamper detection, and reproducibility across independent extraction runs. All results are reported using the full SlideChain Slides Dataset.
@@ 8.3.3 Cost Analysis
On an Ethereum-like L1 environment, per-slide cost ranges from \$0.69 to \$1.23, totaling approximately \$780 for the 1,117-slide dataset. While L1 deployment remains relatively expensive, L2 ecosystems such as Optimism- or Polygon-like rollups reduce these costs by one to two orders of magnitude, making continuous provenance registration feasible for real-world educational pipelines.
@@ 9 Discussion
This work presents SlideChain, a provenance infrastructure that couples multimodal semantic extraction with blockchain-backed verification to support auditability and reproducibility in educational AI systems. Rather than proposing improvements to semantic correctness itself, SlideChain focuses on making the outputs of semantic extraction verifiable, traceable, and comparable over time. Across semantic analysis, blockchain evaluation, and reproducibility experiments, our results highlight both the opportunities and the limitations of current vision–language models when applied to scientific and engineering lecture material, and motivate provenance as a necessary system-level capability.
@@ 10 Conclusion
This paper introduced SlideChain, designed to ensure the integrity, auditability, and long-term reproducibility of multimodal semantic extraction pipelines for educational content. Using the SlideChain Slides Dataset of 1,117 medical imaging lecture slides and four state-of-the-art vision–language models, we presented, to the best of our knowledge, the first empirical study of semantic variability in real-world STEM lecture material. Our results reveal substantial cross-model divergence in both concept extraction and relational grounding, underscoring that modern VLMs—despite their impressive generative capabilities—do not yet provide stable or authoritative interpretations of scientific slides. SlideChain addresses this challenge by anchoring cryptographic commitments of slide-level semantic records to an Ethereum-compatible blockchain. While simple hashing or local logging can detect accidental file changes, they cannot provide immutable timestamps, third-party verifiability, or protection against silent overwriting. In contrast, blockchain-backed commitments establish an append-only, tamper-evident ledger that decouples trust from any single system owner and enables independent auditing of semantic interpretations over time. By storing only compact Keccak-256 hashes on-chain and retaining full semantic records off-chain, SlideChain achieves this integrity guarantee without sacrificing flexibility or scalability. Our evaluations demonstrate that the system exhibits predictable, constant gas usage, near-linear scalability, and stable throughput of approximately one slide per second under batch registration. Tamper-detection experiments achieved 100% accuracy, and dual-run reproducibility tests produced perfect Jaccard similarity across all 4,468 model–slide combinations, confirming deterministic behavior under controlled execution settings. Taken together, these results establish SlideChain as a practical and extensible provenance infrastructure for multimodal educational AI. Rather than attempting to resolve semantic correctness, SlideChain provides a verifiable semantic baseline against which model behavior, drift, and disagreement can be transparently audited. As AI-generated interpretations become increasingly embedded in STEM pedagogy and digital learning ecosystems, such provenance guarantees will be essential for maintaining trust, accountability, and long-term reproducibility.
@@ 11 Future Work
Acknowledgment
We thank Clara Reyes for helpful comments and suggestions that improved the clarity and presentation of the manuscript.
Funding Statement
This research received no external funding.